# Learning Theory for Kernel Bilevel Optimization

### v1

@@ D.2 Maximal inequalities
where $t_{\omega}(z,z^{\prime})\coloneqq\partial_{v}\ell_{out}(\omega,h_{\omega}^{\star}(x),y)\partial_{v}\ell_{out}(\omega,h_{\omega}^{\star}(x^{\prime}),y^{\prime})K(x,x^{\prime})\in\mathcal{T}_{0,1}$. The last term is precisely what Proposition E.4 controls when applying it to the family $\mathcal{T}_{0,1}$ and choosing $\mathbb{D}$ to be $\mathbb{Q}$. Therefore, the following maximal inequality holds by application of Proposition E.4:
@@
where $c(\Omega)$ is a positive constant greater than $1$ the depends only on $\Omega$ and $d$. We obtain a similar inequality for $\partial_{h}\delta_{\omega}^{in}$ by carrying out similar calculations, then applying Proposition E.4 to the family $\mathcal{T}_{1,1}$ and choosing $\mathbb{P}$ for the probability distribution $\mathbb{D}$. The resulting bound is then of the form:
@@
$\mathbb{E}_{\mathbb{P}}{\left[\sup_{\omega\in\Omega}\partial_{\omega,h}^{2}\delta_{\omega}^{in}\right]}^{2}$ $\stackrel{{\scriptstyle(a)}}{{\leq}}\mathbb{E}_{\mathbb{P}}{\left[\sup_{\omega\in\Omega}(\partial_{\omega,h}^{2}\delta_{\omega}^{in})^{2}\right]}$
@@
=$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{l=1}^{d}\mathbb{E}_{\mathbb{P}}{\left[\sup_{\omega\in\Omega}{\left\lVert\mathbb{E}_{(x,y)\sim\mathbb{P}}{\left[\partial_{\omega_{l},v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x),y)K(x,\cdot)\right]}-\frac{1}{n}\sum_{i=1}^{n}\partial_{\omega_{l},v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x_{i}),y_{i})K(x_{i},\cdot)\right\rVert}_{\mathcal{H}}^{2}\right]}$
=$\stackrel{{\scriptstyle(e)}}{{=}}\sum_{l=1}^{d}\mathbb{E}_{\mathbb{P}}{\left[\sup_{\omega\in\Omega}\mathbb{E}_{z,z^{\prime}\sim\mathbb{P}\otimes\mathbb{P}}\left[t_{\omega,l}(z,z^{\prime})\right]+\frac{1}{n^{2}}\sum_{i,j=1}^{n}t_{\omega,l}(z_{i},z_{j})-\frac{2}{n}\sum_{i=1}^{n}\mathbb{E}_{z\sim\mathbb{P}}\left[t_{\omega,l}(z,z_{i})\right]\right]},$
where we introduced $t_{\omega,l}(z,z^{\prime})\coloneqq\partial_{\omega_{l},v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x),y)\partial_{\omega_{l},v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x^{\prime}),y^{\prime})K(x,x^{\prime})\in\mathcal{T}_{2+l,1}$. Here, (a) follows by Cauchy-Schwarz inequality, (b) is obtained by definition of $\partial_{\omega,h}^{2}\delta_{\omega}^{in}$, while (c) uses the general fact that the operator norm of an operator is upper-bounded by its Hilbert-Schmidt norm which is finite in our case by application of Proposition A.2. Moreover, (d) further uses the Hilbert-Schmidt norm of an operator in terms of the norm of its rows, while (e) simply expands the squared RKHS norm and uses the reproducing property in the RKHS $\mathcal{H}$. Each term in the last item (e) is precisely what Proposition E.4 controls when applying it to the families $\mathcal{T}_{2+l,1}$ for $1\leq l\leq d$ and choosing $\mathbb{D}$ to be $\mathbb{P}$. Therefore, the following maximal inequality holds by a direct application of Proposition E.4:
@@
$\stackrel{{\scriptstyle(b)}}{{\coloneqq}}\mathbb{E}_{\mathbb{P}}{\left[\sup_{\omega\in\Omega}{\left\lVert\mathbb{E}_{(x,y)\sim\mathbb{P}}{\left[\partial_{v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x),y)K(x,\cdot)\otimes K(x,\cdot)\right]}-\frac{1}{n}\sum_{i=1}^{n}\partial_{v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x_{i}),y_{i})K(x_{i},\cdot)\otimes K(x_{i},\cdot)\right\rVert}_{\operatorname{op}}^{2}\right]}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\mathbb{E}_{\mathbb{P}}{\left[\sup_{\omega\in\Omega}{\left\lVert\mathbb{E}_{(x,y)\sim\mathbb{P}}{\left[\partial_{v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x),y)K(x,\cdot)\otimes K(x,\cdot)\right]}-\frac{1}{n}\sum_{i=1}^{n}\partial_{v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x_{i}),y_{i})K(x_{i},\cdot)\otimes K(x_{i},\cdot)\right\rVert}_{\operatorname{HS}}^{2}\right]}$
=$\stackrel{{\scriptstyle(d)}}{{=}}\mathbb{E}_{\mathbb{P}}{\left[\sup_{\omega\in\Omega}\mathbb{E}_{z,z^{\prime}\sim\mathbb{P}\otimes\mathbb{P}}\left[t_{\omega}(z,z^{\prime})\right]+\frac{1}{n^{2}}\sum_{i,j=1}^{n}t_{\omega}(z_{i},z_{j})-\frac{2}{n}\sum_{i=1}^{n}\mathbb{E}_{z\sim\mathbb{P}}\left[t_{\omega}(z,z_{i})\right]\right]},$
where we introduced $t_{\omega}(z,z^{\prime})\coloneqq\partial_{v}^{2}\ell_{in}(\omega,x,y)\partial_{v}^{2}\ell_{in}(\omega,x^{\prime},y^{\prime})K^{2}(x,x^{\prime})\in\mathcal{T}_{2,2}$. Here, (a) follows by Cauchy-Schwarz inequality, (b) is obtained by definition of $\partial_{h}^{2}\delta_{\omega}^{in}$, while (c) uses the general fact that the operator norm of an operator is upper-bounded by its Hilbert-Schmidt norm which is finite in our case by application of Proposition A.2. Moreover, (d) further uses the identity in Lemma G.3 for computing the Hilbert-Schmidt norm of sum/expectation of tensor-product operators. The last item (d) is precisely what Proposition E.4 controls when applying it to the family $\mathcal{T}_{2,2}$ and choosing $\mathbb{D}$ to be $\mathbb{P}$. Therefore, the following maximal inequality holds by direct application of Proposition E.4:
@@
D.3 Proof of Theorem 4.1
Theorem D.7 (Generalization bounds).
The following holds under Assumptions (B), (D), (A) and (C):
$\mathbb{E}{\left[\sup_{\omega\in\Omega}{\left\lvert\mathcal{F}(\omega)-\hat{\mathcal{F}}(\omega)\right\rvert}\right]}$ $\lesssim\frac{1}{\lambda m^{\frac{1}{2}}}+\frac{C_{out}}{\lambda^{\frac{5}{4}}n^{\frac{1}{2}}}$
@@
where the constants $C_{in}$ and $C_{out}$ are given in Proposition D.2.
@@
Using the point-wise estimates in Proposition D.4 and taking their supremum over $\Omega$ followed by the expectations over data, the following error bounds hold:
$\mathbb{E}{\left[\sup_{\omega\in\Omega}{\left\lvert\mathcal{F}(\omega)-\hat{\mathcal{F}}(\omega)\right\rvert}\right]}\leq$ $\mathbb{E}_{\mathbb{Q}}{\left[\sup_{\omega\in\Omega}\delta_{\omega}^{out}\right]}+\frac{C_{out}}{\lambda}\mathbb{E}_{\mathbb{P}}{\left[\sup_{\omega\in\Omega}\partial_{h}\delta_{\omega}^{in}\right]},$
@@
Furthermore, we can use the maximal inequalities in Propositions D.5 and D.6 to control each term appearing in the right-hand side of the above inequalities:
$\mathbb{E}{\left[\sup_{\omega\in\Omega}{\left\lvert\mathcal{F}(\omega)-\hat{\mathcal{F}}(\omega)\right\rvert}\right]}\leq$ $R{\left(m^{-\frac{1}{2}}\lambda^{-1}+C_{out}n^{-\frac{1}{2}}\lambda^{-(1+\frac{1}{4})}\right)}$
$\mathbb{E}{\left[\sup_{\omega\in\Omega}{\left\lVert\nabla\mathcal{F}(\omega)-\widehat{\mathcal{F}}(\omega)\right\rVert}\right]}\leq$ $R\Big{(}m^{-\frac{1}{2}}\lambda^{-1}d^{\frac{1}{2}}+C_{in}m^{-\frac{1}{2}}\lambda^{-(1+\frac{1}{4})}+C_{out}n^{-\frac{1}{2}}\lambda^{-(1+\frac{1}{4})}{\left(1+2\frac{C_{in}}{\lambda}+\frac{C_{in}^{2}}{\lambda^{2}}\right)}$
@@
where the constant $R$ depends only on the Lipschitz constants $\operatorname{Lip}_{in}$, $\operatorname{Lip}_{out}$, the upper-bounds $M_{in}$, $M_{out}$, the bound $\kappa$ on the kernel, the set $\Omega$ and the dimension $d$. Rearranging the obtained upper-bounds concludes the proof. ∎
Appendix E Maximal Inequalities for Bounded and Lipschitz Family of Functions

### v3

@@ E.2 Maximal inequalities
where $t_{\omega}(z,z^{\prime})\coloneqq\partial_{v}\ell_{out}(\omega,h_{\omega}^{\star}(x),y)\partial_{v}\ell_{out}(\omega,h_{\omega}^{\star}(x^{\prime}),y^{\prime})K(x,x^{\prime})\in\mathcal{T}_{0,1}$. The last term is precisely what Proposition˜F.4 controls when applying it to the family $\mathcal{T}_{0,1}$ and choosing $\mathbb{D}$ to be $\mathbb{Q}$. Therefore, the following maximal inequality holds by application of Proposition˜F.4:
@@
where $c(\Omega)$ is a positive constant greater than $1$ that depends only on $\Omega$ and $d$. We obtain a similar inequality for $\partial_{h}\delta_{\omega}^{in}$ by carrying out similar calculations, then applying Proposition˜F.4 to the family $\mathcal{T}_{1,1}$ and choosing $\mathbb{P}$ for the probability distribution $\mathbb{D}$. The resulting bound is then of the form:
@@
$\mathbb{E}_{\mathbb{P}}\left[\sup_{\omega\in\Omega}\partial_{\omega,h}^{2}\delta_{\omega}^{in}\right]^{2}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $\mathbb{E}_{\mathbb{P}}\left[\sup_{\omega\in\Omega}(\partial_{\omega,h}^{2}\delta_{\omega}^{in})^{2}\right]$
@@
$\stackrel{{\scriptstyle(d)}}{{=}}$ $\sum_{l=1}^{d}\mathbb{E}_{\mathbb{P}}\left[\sup_{\omega\in\Omega}\left\lVert\mathbb{E}_{(x,y)\sim\mathbb{P}}\left[\partial_{\omega_{l},v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x),y)K(x,\cdot)\right]-\frac{1}{n}\sum_{i=1}^{n}\partial_{\omega_{l},v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x_{i}),y_{i})K(x_{i},\cdot)\right\rVert_{\mathcal{H}}^{2}\right]$
$\stackrel{{\scriptstyle(e)}}{{=}}$ $\sum_{l=1}^{d}\mathbb{E}_{\mathbb{P}}\left[\sup_{\omega\in\Omega}\mathbb{E}_{z,z^{\prime}\sim\mathbb{P}\otimes\mathbb{P}}\left[t_{\omega,l}(z,z^{\prime})\right]+\frac{1}{n^{2}}\sum_{i,j=1}^{n}t_{\omega,l}(z_{i},z_{j})-\frac{2}{n}\sum_{i=1}^{n}\mathbb{E}_{z\sim\mathbb{P}}\left[t_{\omega,l}(z,z_{i})\right]\right],$
where we introduced $t_{\omega,l}(z,z^{\prime})\coloneqq\partial_{\omega_{l},v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x),y)\partial_{\omega_{l},v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x^{\prime}),y^{\prime})K(x,x^{\prime})\in\mathcal{T}_{2+l,1}$. Here, (a) follows from the Cauchy-Schwarz inequality, (b) is obtained by definition of $\partial_{\omega,h}^{2}\delta_{\omega}^{in}$, while (c) uses the general fact that the operator norm of an operator is upper-bounded by its Hilbert-Schmidt norm which is finite in our case by application of Proposition˜B.2. Moreover, (d) further uses the Hilbert-Schmidt norm of an operator in terms of the norm of its rows, while (e) simply expands the squared RKHS norm and uses the reproducing property in the RKHS $\mathcal{H}$. Each term in the last item (e) is precisely what Proposition˜F.4 controls when applying it to the families $\mathcal{T}_{2+l,1}$ for $1\leq l\leq d$ and choosing $\mathbb{D}$ to be $\mathbb{P}$. Therefore, the following maximal inequality holds by a direct application of Proposition˜F.4:
@@
$\stackrel{{\scriptstyle(b)}}{{\coloneqq}}$ $\mathbb{E}_{\mathbb{P}}\Bigg[\sup_{\omega\in\Omega}\Bigg\|\mathbb{E}_{(x,y)\sim\mathbb{P}}\left[\partial_{v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x),y)K(x,\cdot)\otimes K(x,\cdot)\right]$
$\qquad\qquad\qquad\qquad-\frac{1}{n}\sum_{i=1}^{n}\partial_{v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x_{i}),y_{i})K(x_{i},\cdot)\otimes K(x_{i},\cdot)\Bigg\|_{\operatorname{op}}^{2}\Bigg]$
$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $\mathbb{E}_{\mathbb{P}}\Bigg[\sup_{\omega\in\Omega}\Bigg\|\mathbb{E}_{(x,y)\sim\mathbb{P}}\left[\partial_{v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x),y)K(x,\cdot)\otimes K(x,\cdot)\right]$
$\qquad\qquad\qquad\qquad-\frac{1}{n}\sum_{i=1}^{n}\partial_{v}^{2}\ell_{in}(\omega,h_{\omega}^{\star}(x_{i}),y_{i})K(x_{i},\cdot)\otimes K(x_{i},\cdot)\Bigg\|_{\operatorname{HS}}^{2}\Bigg]$
$\stackrel{{\scriptstyle(d)}}{{=}}$ $\mathbb{E}_{\mathbb{P}}\left[\sup_{\omega\in\Omega}\mathbb{E}_{z,z^{\prime}\sim\mathbb{P}\otimes\mathbb{P}}\left[t_{\omega}(z,z^{\prime})\right]+\frac{1}{n^{2}}\sum_{i,j=1}^{n}t_{\omega}(z_{i},z_{j})-\frac{2}{n}\sum_{i=1}^{n}\mathbb{E}_{z\sim\mathbb{P}}\left[t_{\omega}(z,z_{i})\right]\right],$
where we introduced $t_{\omega}(z,z^{\prime})\coloneqq\partial_{v}^{2}\ell_{in}(\omega,x,y)\partial_{v}^{2}\ell_{in}(\omega,x^{\prime},y^{\prime})K^{2}(x,x^{\prime})\in\mathcal{T}_{2,2}$. Here, (a) follows from the Cauchy-Schwarz inequality, (b) is obtained by definition of $\partial_{h}^{2}\delta_{\omega}^{in}$, while (c) uses the general fact that the operator norm of an operator is upper-bounded by its Hilbert-Schmidt norm which is finite in our case by application of Proposition˜B.2. Moreover, (d) further uses the identity in Lemma˜H.3 for computing the Hilbert-Schmidt norm of sum/expectation of tensor-product operators. The last item (d) is precisely what Proposition˜F.4 controls when applying it to the family $\mathcal{T}_{2,2}$ and choosing $\mathbb{D}$ to be $\mathbb{P}$. Therefore, the following maximal inequality holds by direct application of Proposition˜F.4:
@@
E.3 Proof of Theorem˜4.1
Theorem E.7 (Generalization bounds).
The following holds under Assumptions˜(A), (C), (E), (B) and (D):
$\mathbb{E}\left[\sup_{\omega\in\Omega}\left\lvert\mathcal{F}(\omega)-\widehat{\mathcal{F}}(\omega)\right\rvert\right$] $\lesssim\frac{1}{\lambda m^{\frac{1}{2}}}+\frac{C_{out}}{\lambda^{\frac{5}{4}}n^{\frac{1}{2}}},$
@@
where the constants $C_{in}$ and $C_{out}$ are given in Proposition˜E.2.
@@
Using the point-wise estimates in Proposition˜E.4 and taking their supremum over $\Omega$ followed by the expectations over data, the following error bounds hold:
$\mathbb{E}\left[\sup_{\omega\in\Omega}\left\lvert\mathcal{F}(\omega)-\widehat{\mathcal{F}}(\omega)\right\rvert\right$$\leq$ $\mathbb{E}_{\mathbb{Q}}\left[\sup_{\omega\in\Omega}\delta_{\omega}^{out}\right]+\frac{C_{out}}{\lambda}\mathbb{E}_{\mathbb{P}}\left[\sup_{\omega\in\Omega}\partial_{h}\delta_{\omega}^{in}\right],$
@@
Furthermore, we can use the maximal inequalities in Propositions˜E.5 and E.6 to control each term appearing in the right-hand side of the above inequalities:
$\mathbb{E}\left[\sup_{\omega\in\Omega}\left\lvert\mathcal{F}(\omega)-\widehat{\mathcal{F}}(\omega)\right\rvert\right$$\leq$ $R\left(m^{-\frac{1}{2}}\lambda^{-1}+C_{out}n^{-\frac{1}{2}}\lambda^{-(1+\frac{1}{4})}\right),$
$\mathbb{E}\left[\sup_{\omega\in\Omega}\left\lVert\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\right\rVert\right]\leq$ $R\Big(m^{-\frac{1}{2}}\lambda^{-1}d^{\frac{1}{2}}+C_{in}m^{-\frac{1}{2}}\lambda^{-(1+\frac{1}{4})}$
$\qquad+C_{out}n^{-\frac{1}{2}}\lambda^{-(1+\frac{1}{4})}\left(1+2\frac{C_{in}}{\lambda}+\frac{C_{in}^{2}}{\lambda^{2}}\right)$
@@
where the constant $R$ depends only on the Lipschitz constants $\operatorname{Lip}_{in}$ and $\operatorname{Lip}_{out}$, the upper bounds $M_{in}$ and $M_{out}$, the bound $\kappa$ on the kernel, the set $\Omega$, and the dimension $d$. Rearranging the obtained upper bounds concludes the proof. ∎
E.4 Generalization for bilevel gradient methods
Proof of Corollary˜4.2.
Consider that $\inf_{\omega,v,y}\ell_{out}(\omega,v,y)-c\|\omega\|^{2}\geq 0$, which entails $\ell_{out}(\omega,v,y)\geq c\|\omega\|^{2}$ for all $v,y$. Using Proposition˜D.1 and setting $B=\sup_{y\in\mathcal{Y}}|\partial_{v}\ell_{in}(\omega_{0},0,y)|$, we have almost surely that:
$\widehat{\mathcal{F}}(\omega_{0})\leq\max_{|v|\leq\frac{B\kappa}{\lambda},y\in\mathcal{Y}}\ell_{out}(\omega_{0},v,y)\eqqcolon\bar{\ell}.$
Therefore, for any $\omega$ such that $\widehat{\mathcal{F}}(\omega)\leq\widehat{\mathcal{F}}(\omega_{0})$, we have $\|\omega\|^{2}\leq\bar{\ell}/c$. Define $\Omega$ as the ball of radius $\sqrt{\bar{\ell}/c}$ centered at 0. Using the fact that $\widehat{\nabla\mathcal{F}}$$\nabla\widehat{\mathcal{F}}$ in Proposition˜3.1 and the representation in Equation˜4, it is clear from Proposition˜D.2 and Assumption˜(D) that $\nabla\widehat{\mathcal{F}}$ is Lipschitz on $\Omega$ with a deterministic constant $L$. It follows from standard results on gradient descent for nonconvex $\widehat{\mathcal{F}}$ with Lipschitz gradient (see, e.g., [beck2014introduction, Theorems 4.25, 4.26]) that if we take $\bar{\eta}=1/L$, then almost surely:
  • $\widehat{\mathcal{F}}(\omega_{t})\leq\widehat{\mathcal{F}}(\omega_{0})$ and $\omega_{t}\in\Omega$ for all $t\geq 0$.
  • $\nabla\widehat{\mathcal{F}}(\omega_{t})\to 0$ as $t\to\infty$.
The corollary then follows by combining Proposition˜3.1 and the uniform bound in Theorem˜4.1. ∎
Bilevel projected gradient descent. Considering the constrained (KBO) problem and assuming that $\mathcal{C}$ is convex and compact, the projected gradient descent initialized at $\omega_{0}\in\mathcal{C}$ iterates the following recursion $\omega_{t+1}=\Pi_{\mathcal{C}}(\omega_{t}-\eta\nabla\widehat{\mathcal{F}}(\omega_{t}))$ for all $t\geq 0$, where $\Pi_{\mathcal{C}}$ denotes the orthogonal projection onto $\mathcal{C}$ and $\eta>0$ is the step size. The algorithmic requirements are the same as the gradient descent algorithm, with the additional cost of computing the projection, which is typically cheap for basic sets such as balls. In the constrained setting, the optimality condition should take the constraints into account. To this end, we consider the gradient mappings $\widehat{G}_{\eta}\colon\omega\mapsto\frac{1}{\eta}(\omega-\Pi_{\mathcal{C}}(\omega-\eta\nabla\widehat{\mathcal{F}}(\omega)))$ and $G_{\eta}\colon\omega\mapsto\frac{1}{\eta}\left(\omega-\Pi_{\mathcal{C}}(\omega-\eta\nabla\mathcal{F}(\omega))\right)$ [beck2017first, Section 10.3]. This captures the stationarity of the recursion, and any local minimum of $\mathcal{F}$ on $\mathcal{C}$ satisfies $G_{\eta}=0$ for all $\eta>0$.
Corollary E.8 (Generalization for bilevel projected gradient descent).
Consider Assumptions˜(A), (C), (E), (B) and (D) and fix $\lambda>0$. Assume further that $\mathbf{K}$ in ($\widehat{\text{KBO}}$) is almost surely definite, and that $\mathcal{C}$ is convex and compact. Fix $\omega_{0}\in\mathcal{C}$ and let $\omega_{t+1}=\Pi_{\mathcal{C}}(\omega_{t}-\eta\nabla\widehat{\mathcal{F}}(\omega_{t}))$, where $\eta>0$ is the step size and $t\geq 0$ is the iteration index. Then, there exist constants $\bar{\eta}>0$ and $\bar{c}>0$ such that for any $0<\eta<\bar{\eta}$ and $t>0$, the following holds:
$\mathbb{E}\Big[\min_{i=0,\ldots,t}\left\|G_{\eta}(\omega_{i})\right\|\Big]\leq\bar{c}\left(\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{n}}+\frac{1}{\sqrt{t+1}}\right),\mathbb{E}\Big[\underset{i\to\infty}{\lim\sup}\left\|G_{\eta}(\omega_{i})\right\|\Big]\leq\bar{c}\left(\frac{1}{\sqrt{m}}+\frac{1}{\sqrt{n}}\right).$
Proof.
We choose $\Omega=\mathcal{C}$. All iterates obviously remain in $\Omega$. Similarly as in the proof of Corollary˜4.2, we know that $\nabla\widehat{\mathcal{F}}$ is Lipschitz and that $\mathcal{F}$ is bounded on $\Omega$ with deterministic constants. It then follows from classical analysis on the nonconvex projected gradient algorithm (see, e.g., [beck2017first, Theorem 10.15]), that for sufficiently small $\eta$, we have almost surely that $\min_{i=0,\ldots,t}\left\|\widehat{G}_{\eta}(\omega_{i})\right\|\leq\bar{c}/\sqrt{t+1}$ for a deterministic constant $\bar{c}>0$, and that $\widehat{G}_{\eta}(\omega_{i})\to 0$ as $i\to\infty$. Using the fact that the orthogonal projection is $1$-Lipschitz, (see, e.g., [beck2017first, Theorem 6.42]), we also have for all $\omega\in\Omega$:
The result follows by combining Proposition˜3.1 and the uniform bound in Theorem˜4.1. ∎
Appendix F Maximal Inequalities for Bounded and Lipschitz Family of Functions